# Limit cycles and invariant algebraic curves

### v1

@@ 1. Introduction and statement of the main results
For an account of known results on planar polynomial vector fields with algebraic invariant curves see [7, Chapter $8$] and its references. The interest on this subject goes back to the seminal works of Darboux and Poincaré.
@@
We observe that (2) can be seen as a generalization of [9, Theorem $1$], where it was proved that $\mathcal{H}(n+1)\geqslant\mathcal{H}(n)+1$. Nevertheless notice that this last inequality is stronger than (2) when $m=1,$ as $\operatorname{Har}(1)=0.$
@@
Let $\mathcal{H}_{K}(n)$ denote the maximum number of limit cycles of (3) in the first quadrant, where $p$ and $q$ vary among all polynomials of degree at most n1n-1. Next table gives their exact values for $n=1,2$ and some lower bounds for $n\geqslant 3$.
@@
$\mathcal{H}_{K}(n)$ 0 0 66 $13$ $22$ $28$ $37$ $53$ $O(n^{2}\ln n)$
Case $n=1$ is trivial, as linear systems do not have limit cycles. When $2\leqslant n\leqslant 5$, it follows from previous works, see the references in the proof. In particular, for $3\leqslant n\leqslant 5$ the values have been obtained by high order Andronov–Hopf bifurcations. The results for bigger $n$ are new and follow from our approach based on Theorem A.
@@
Let $\mathcal{H}_{\scriptscriptstyle{\square}}(n)$ denote the maximum number of limit cycles of (4) in the open square $\mathcal{Q}$, where $p$ and $q$ vary among all polynomials of degree at most n2n-2. Next table gives their exact values for $n=2,3$ and some lower bounds for $n\geqslant 4$.
@@
$\mathcal{H}_{\scriptscriptstyle{\square}}(n)$ 0 11 55 55 55 $13$ $28$ $37$ $53$ $O(n^{2}\ln n)$
As in Corollary 4, the proof for $n=2$ is easy: it is an integrable system. Cases $3\leqslant n\leqslant 4$ follow from previous results. Again the main tool when $n=4$ is the study of a high order Andronov–Hopf bifurcation. The results for $n=5$ and $6$ follow trivially from the one for $n=4,$ and for bigger $n$ they are new and a consequence of the main result of this work. It is clear that the cases $n=5$ and $6$ deserve more attention. It is sure that a careful study of them will give higher lower bounds.
@@ 2. Preliminaries
The first result that we will use is a simplified version of [9, Theorem $2$].
@@
It is clear that $C=0$ is an invariant algebraic curve for (5) because $C_{x}P+C_{y}Q=\big(\alpha C_{x}+\beta C_{y}\big)C.$ Moreover, clearly $Y\in\Sigma_{C}^{c}$. Let $\gamma$ be an oval of $C=0$, i.e. a bounded connected component of $C=0$ without singular points. Since $D=0$ does not intersect $\gamma$ and $\nabla C$ never vanishes at it, we have that $Y$ does not have singularities at $\gamma$ and thus it is a periodic orbit of $Y$. It can be parameterized by time and let $T$ denote its minimal period. To prove that it is a hyperbolic limit cycle it suffices to show that next integral does not vanishes, see [7, Section 7.3]. We have,
@@ 3. Proofs
Suppose first $\mathcal{H}(n)<\infty$. It follows from Theorem 1(a)(a) that there is $X=(U,V)$ such that $\pi(X)=\mathcal{H}(n)$ and all its limit cycles are hyperbolic. Let $B\subset\mathbb{R}^{2}$ be a closed ball containing all the limit cycles of $X$ in its interior. Translating and scaling $X$ if necessary, we can suppose $B\cap C^{-1}(\{0\})=\emptyset$.
@@
It follows from [12, Theorem $2.1$] that there is $M>0$ such that for each $n\geqslant 3$, $\mathcal{H}(n)\geqslant(n^{2}\log_{2}n)/2-Mn^{2}\log_{2}\log_{2}n.$ Hence from (6) we have that, for $n$ big enough, $\mathcal{H}_{C}(n)$ increases at least as fast as $O(n^{2}\ln n)$. ∎
@@
We have from [1] and [17] that $\mathcal{H}_{K}(2)=0$ and $\mathcal{H}_{K}(3)\geqslant 6$. From [3] we know that $\mathcal{H}_{K}(4)\geqslant 13$ and $\mathcal{H}_{K}(5)\geqslant 22$.
From Theorem A we have $\mathcal{H}_{K}(n+2)\geqslant\mathcal{H}(n)$ and observe that from the proof of this theorem we can choose these limit cycles to be at the first quadrant. Hence the corollary follows by using the known lower bounds of $\mathcal{H}(n)$ given in Table 1.
Finally we observe that the vector fields provided by Theorem A are non-degenerated, i.e. they have at most finitely many singularities at $xy=0$. ∎
@@
I We know from [1] and [14, Theorem $2.5.4$] that $\mathcal{H}_{\scriptscriptstyle{\square}}(2)=0$ and $\mathcal{H}_{\scriptscriptstyle{\square}}(3)=1$. From [8, 10] we have $\mathcal{H}_{\scriptscriptstyle{\square}}(4)\geqslant 5$. Trivially, $\mathcal{H}_{\scriptscriptstyle{\square}}(5)$ and $\mathcal{H}_{\scriptscriptstyle{\square}}(6)$ are at least $5$. From Theorem A we know that $\mathcal{H}_{\scriptscriptstyle{\square}}(n+4)\geqslant\mathcal{H}(n)$ and, as in the previous corollary, the results for $n\geqslant 7$ follow from the known lower bounds for $\mathcal{H}(n)$.
Moreover, all these limit cycles are hyperbolic and can be located to be realizable by the game, that is, in the open square $\mathcal{Q}.$

### v2

@@ 1. Introduction and statement of the main results
For an account of known results on planar polynomial vector fields with algebraic invariant curves see [8, Chapter $8$] and its references. The interest on this subject goes back to the seminal works of Darboux and Poincaré.
@@
We observe that (2) can be seen as a generalization of [10, Theorem $1$], where it was proved that $\mathcal{H}(n+1)\geqslant\mathcal{H}(n)+1$. Nevertheless notice that this last inequality is stronger than (2) when $m=1,$ as $\operatorname{Har}(1)=0.$
@@
Since $\deg C=2$ and $xy=0$ has no ovals, by Theorem A, $\mathcal{H}_{K}(n)\geqslant\mathcal{H}(n-2).$ In a first version of this work we had used this inequality for $n\geqslant 5.$ In a recent 2025 meeting the authors were communicated by Plesa that the technique employed by him in [18] could improve that lower bound to $\mathcal{H}_{K}(n)\geqslant\mathcal{H}(n-1)$. We have used this technique in next two results. We greatly appreciate his suggestion. As we will see in the proof of next corollaries, this improvement can be done due to the very particular and simple shape of the invariant algebraic curves in these cases.
@@
Let $\mathcal{H}_{K}(n)$ denote the maximum number of limit cycles of (3) in the first quadrant, where $p$ and $q$ vary among all polynomials of degree at most n1n-1. Then $\mathcal{H}_{K}(n)\geqslant\mathcal{H}(n-1)$ and next table gives their exact values for $n=1,2$ and some lower bounds for $n\geqslant 3$.
@@
$\mathcal{H}_{K}(n)$ 0 0 66 $13$ $28$ $37$ $53$ $74$ $O(n^{2}\ln n)$
In case $n=1$ the result is simply because linear systems do not have limit cycles. When $n\in\{2,3\}$, it follows from previous works, see the references in the proof. For the case $n=3$ the value was obtained by a high order Andronov–Hopf bifurcation. By using this last method it is proved in [3] that $\mathcal{H}_{K}(4)\geqslant 13$ and $\mathcal{H}_{K}(5)\geqslant 22.$ Our approach recovers the same lower bound when $n=4$ and improves it when $n=5.$ As far as we know, the results for $n\geqslant 6$ and the asymptotic lower bound for large $n$ are new. Moreover, from our proof we also obtain that all existing configurations of hyperbolic limit cycles for polynomial systems of degree n1n-1 are realizable in the first quadrant for Kolmogorov systems of degree $n.$ From this point of view the above two results showing that $\mathcal{H}_{K}(4)\geqslant 13$ are different. While the 13 limit cycles given in [3] are nested, the 13 given by our approach are not nested. Indeed, until now only 12 nested limit cycles have been found for cubic systems.
@@
Let $\mathcal{H}_{\scriptscriptstyle{\square}}(n)$ denote the maximum number of limit cycles of (4) in the open square $\mathcal{Q}$, where $p$ and $q$ vary among all polynomials of degree at most n2n-2. Then $\mathcal{H}_{\scriptscriptstyle{\square}}(n)\geqslant\mathcal{H}(n-2)$ and next table gives their exact values for $n=2,3$ and some lower bounds for $n\geqslant 4$.
@@
$\mathcal{H}_{\scriptscriptstyle{\square}}(n)$ 0 11 55 $13$ $28$ $37$ $53$ $74$ $96$ $O(n^{2}\ln n)$
Cases $2\leqslant n\leqslant 4$ follow from previous known results, see the references in the proof. Again the main tool when $n=4$ was the study of a high order Andronov–Hopf bifurcation. For $n\geqslant 5,$ as well as for the asymptotic lower bound for big $n$, the results are new. Similarly that in the Kolmogorov case, all existing configurations of hyperbolic limit cycles for polynomial systems of degree n2n-2 are realizable in $\mathcal{Q}$ for systems of degree $n$ of the form (4).
@@ 2. Preliminaries
The first result that we will use is a simplified version of [10, Theorem $2$].
@@
It is clear that $C=0$ is an invariant algebraic curve for (5) because $C_{x}P+C_{y}Q=\big(\alpha C_{x}+\beta C_{y}\big)C.$ Moreover, clearly $Y\in\Sigma_{C}^{c}$. Let $\gamma$ be an oval of $C=0$, i.e. a bounded connected component of $C=0$ without singular points. Since $D=0$ does not intersect $\gamma$ and $\nabla C$ never vanishes at it, we have that $Y$ does not have singularities at $\gamma$ and thus it is a periodic orbit of $Y$. It can be parameterized by time and let $T$ denote its minimal period. To prove that it is a hyperbolic limit cycle it suffices to show that next integral does not vanishes, see [8, Section 7.3]. We have,
@@ 3. Proofs
Suppose first $\mathcal{H}(n)<\infty$. It follows from Theorem 1(a)(a) that there is $X=(U,V)$ such that $\pi(X)=\mathcal{H}(n)$ and all its limit cycles are hyperbolic. Let $B\subset\mathbb{R}^{2}$ be a closed ball containing all the limit cycles introduced so far (i.e. all the limit cycles of $X$) in its interior. Translating and scaling $X$ if necessary, we can suppose $B\cap C^{-1}(\{0\})=\emptyset$.
@@
It follows from [13, Theorem $2.1$] that there is $M>0$ such that for each $n\geqslant 3$, $\mathcal{H}(n)\geqslant(n^{2}\log_{2}n)/2-Mn^{2}\log_{2}\log_{2}n.$ Hence from (6) we have that, for $n$ big enough, $\mathcal{H}_{C}(n)$ increases at least as fast as $O(n^{2}\ln n)$. ∎
@@
Case $n=1$ follows from the fact that linear systems do not have limit cycles. For $n=2$ it was proved by Bautin [1] that quadratic Kolmogorov vector fields cannot have limit cycles. See also [7]*Section $5$. By using [17] we get $\mathcal{H}_{K}(3)\geqslant 6$. From Theorem A we obtain the asymptotic expansion for $n$ large.
To end the proof it suffices to prove $\mathcal{H}_{K}(n)\geqslant\mathcal{H}(n-1)$, $n\geqslant 1,$ and use the known lower bounds of $\mathcal{H}(n-1)$ for $n\geqslant 5$ given in Table 1. We adapt to our setting the technique developed by Plesa  [18]. To this end, similarly to the proof of Theorem A suppose $\mathcal{H}(n-1)<\infty$ and consider a vector field $X=(U,V)$ of degree n1n-1 such that $\pi(X)=\mathcal{H}(n-1)$ and all its limit cycles hyperbolic, which can be supposed by Theorem 1. Let $B\subset\mathbb{R}^{2}$ be a closed ball containing all the limit cycles of $X$ in its interior. Consider the family of vector fields of degree $n,$ $X_{\mu}=(R_{\mu},S_{\mu}),$ given by
$R_{\mu}=(1+\mu x)U,\quad S_{\mu}=(1+\mu y)V,$
where $\mu$ is a non-negative parameter. Notice that $X_{0}=X,$ and the following important properties of $X_{\mu},$ when $\mu>0$ is small enough:
  • It has the invariant straight lines $x=-1/\mu$, and $y=-1/\mu$. Moreover, since $B$ is compact, such lines do not intersect $B.$
  • Restricted to $B,$ and again due to its compactness, $X_{\mu}$ can be seen as an arbitrary small perturbation of $X_{0}=X.$ This is no more true far away from $B.$
As a consequence, since all the limit cycles of $X$ are hyperbolic and they are contained in $B$ it follows that if $\mu>0$ is small enough, all them persist for $X_{\mu}.$ Finally, if we consider the new coordinates $x_{1}=1+\mu x$, $y_{1}=1+\mu y$, then $X_{\mu}$ becomes a Kolmogorov vector field of the form (3), of degree $n$ and with at least $\mathcal{H}(n-1)$ hyperbolic limit cycles in the first quadrant and keep their configuration. The case $\mathcal{H}(n-1)=\infty$ follows similarly to Theorem A. ∎
@@
When $n=2$ the system is easily integrable and $\mathcal{H}_{\scriptscriptstyle{\square}}(2)=0$. Case $n=3$ follows from [14, Theorem $2.5.4$]. From [9, 11] we obtain $\mathcal{H}_{\scriptscriptstyle{\square}}(4)\geqslant 5$. As in Corollary 3, from Theorem A we get the asymptotic expansion for $n$ large.
To end the proof it suffices to prove that $\mathcal{H}_{\scriptscriptstyle{\square}}(n)\geqslant\mathcal{H}(n-2)$, $n\geqslant 2,$ and use the lower bounds of $\mathcal{H}(n-1)$ for $n\geqslant 5$ of Table 1. The proof of this inequality follows by adapting the proof of the similar one used in Corollary 3. We only point out some minor differences. In this case, when $\mathcal{H}(n-2)<\infty$ we start with a vector field $X=(U,V)$ of degree n2n-2 such that $\pi(X)=\mathcal{H}(n-2)$ and all its limit cycles hyperbolic, by Theorem 1. Then, the suitable 1-parameter family of vector fields $X_{\mu}=(R_{\mu},S_{\mu})$ of degree $n$ that allows us to prove the result is
$R_{\mu}=(1+\mu x)(1-\mu x)U,\quad S_{\mu}=(1+\mu y)(1-\mu y)V.$
The case $\mathcal{H}(n-2)=\infty$ also follows similarly. We skip the details. ∎